MBER SPECTROSCOPY OF ISOCYANIC ACID (HNCO)

\underline{BRIAN MEEHAN}, and J. S. MUENTER, Department of Chemistry, University of Rochester, Rochester, NY 14627.

Recent studies of vibrationally mediated photodissociation of HNCO ^{a} point to the need for better understanding of the vibrational dependence of HNCO electronic properties. A necessary preliminary to excited state measurements is accurate characterization of ground state properties. The present work describes molecular beam electric resonance studies of several rotational levels in the ground vibrational state of isocyanic acid. Dipole moment and hyperfine results will be presented.

aS. S. Brown, H. L. Berghout, and F. F. Crim, J. Chem. Phys. 102, 8440 (1995).